### v2

# Exploring the Impact of Dataset Statistical Effect Size on Model Performance and Data Sample Size Sufficiency

Abstract
Having a sufficient quantity of quality data is a critical enabler of training effective machine learning models. Being able to effectively determine the adequacy of a dataset prior to training and evaluating a model’s performance would be an essential tool for anyone engaged in experimental design or data collection. However, despite the need for it, the ability to prospectively assess data sufficiency remains an elusive capability. We report here on two experiments undertaken in an attempt to better ascertain whether or not basic descriptive statistical measures can be indicative of how effective a dataset will be at training a resulting model. Leveraging the effect size of our features, this work first explores whether or not a correlation exists between effect size, and resulting model performance (theorizing that the magnitude of the distinction between classes could correlate to a classifier’s resulting success). We then explore whether or not the magnitude of the effect size will impact the rate of convergence of our learning rate, (theorizing again that a greater effect size may indicate that the model will converge more rapidly, and with a smaller sample size needed). Our results appear to indicate that this is not an effective heuristic for determining adequate sample size or projecting model performance, and therefore that additional work is still needed to better prospectively assess adequacy of data.
Codehttps://colab.research.google.com/drive/17cZ8pWYAQzCcENLgoUS01cRj5uLWi8ta?usp=sharing
Datasetshttp://archive.ics.uci.edu/dataset/2/adult
Introduction
The sufficiency of a dataset, both in terms of size and representativeness, is critical to training an effective Machine Learning model. Failure to train on data of adequate quality is likely to result in models that dramatically under-perform in production environments (Friedland 2024).
Training machine learning models on inadequate data presents several challenges that can significantly impact model performance and generalizability. While the problem of ’inadequacy’ of data is a multifaceted one, one significant component of it is a lack of sufficient data volume, which in-turn likely impacts the representativeness of the data sample towards the overall population undermining the ability of a model to learn meaningful patterns, resulting in overfitting, where the model performs well on the training data but struggles to generalize to unseen data leading to poor outcomes when deployed in real-world scenarios. This owes to the fact that if the dataset used for training does not accurately reflect the conditions and distributions the model will encounter in real-world applications, the model will struggle to generalize beyond the narrow scope of its training (Masiha et al. 2021). Consequently, the model’s assumptions about the domain may be incorrect, leading to suboptimal decision-making in practical applications.
In cases where the dataset is particularly small or inadequate, machine learning models—especially complex ones like deep learning architectures—fail to learn the intricate relationships within the data. Complex models require large amounts of data to effectively capture subtle patterns (Keshari et al. 2020). Without sufficient data, the model may not converge, or it may underfit, meaning it performs poorly on both the training and test sets due to an inability to learn the underlying patterns.
Moreover, the use of inadequate data increases variability in model performance. Small datasets are more sensitive to changes in initialization values, training procedures, or even small differences in the data samples used during training (Dodge et al. 2020). As a result, models trained on inadequate data exhibit inconsistent and unpredictable behavior, making them unreliable in different testing or production environments. Issues such as overfitting, bias toward dominant classes, poor handling of noisy data, and a lack of generalizability can result from insufficient, non-representative, or poor-quality data. Ensuring access to high-quality, diverse, and sufficiently large datasets is essential for developing models that are both robust and reliable in their predictions.
Having a reliable means of determining the quantity of training data required to effectively train a model would be an incredibly useful tool for researchers to have. This owes both to the upfront challenges and associated costs of data collection, and with the advent of increasingly costly training runs, the investment costs associated with training.
This is a widely recognized need across research domains, with many experimental protocols requiring power analyses, or equivalent statistical studies to prospectively determine the requisite amount of data collection required to run studies effectively.(Bausell and Li 2002)
However, in spite of the obvious utility of such an assessment, it remains elusive for machine learning model developers. As already noted, many methods for external dataset-based validation but are post-hoc model dependent.
Background
Prospective Analysis of Data Sufficiency
In most academic disciplines, conducting a sufficiently robust study design is arguably more important than the statistical analysis that succeeds it. After-all, a poorly designed study may be unsalvageable after the fact, whereas a poorly analyzed study can simply be re-analyzed once errors in methodology are determined (BMJ 2020).
A critical component in study design is the determination of the appropriate sample size. The sample size must be large enough to establish statistical significance of any potential findings, yet not so large as to unnecessarily burden researchers with the (often costly) acquisition of data.
Attempting to determine data sufficiency for machine learning (ML) models is a particularly vexing challenge given the nature of machine learning. Unlike statistics, ML does not attempt to asset any factual relationship between label and feature set (Bennett et al. 2022); Rather ML models are almost mercenary in nature. Charged with best accomplishing a given prediction or classification task, irrespective of any actual underlying statistical relationship.
This relaxed focus allows for a more expansive inclusion of features for whom a statistically significant relationship to the label may not exist.
Current approaches to determining sufficient sample size
There is a lack of approaches that exist for determining a sufficient sample size to train ML models. The absence of a robust pre-hoc approach that is broadly applicable remains a significant challenge in machine learning, largely stemming from the limited research conducted in this area (Balki et al. 2019).
The most common current approach is empirical analysis. This post-hoc method occurs when the sample size for training is incrementally increased and the learning curve is analyzed to see where the point of diminishing returns occurs in relation to sample size. This enables the identification of the correlation between sample size and model performance.
The only existing approaches that can be considered a pre-hoc approach are model-based. These approaches utilize the parameters of an algorithm to determine sample size. Some models, like neural networks, typically need more data to find a pattern and develop a function that maximizes accuracy when compared to simpler models like decision trees. This is due to the nature of how the model is developed to predict. The more parameters a model has, the more complex it tends to be. Neural networks typically have multiple parameters in each of the hidden layers whereas other algorithms only have one layer. There are mathematical equations like those of Baum and Hausler that allow one to determine a sufficient number of samples through this model-based framework. However, this framework is not optimal for all tasks and has its trade-offs.(Balki et al. 2019)
Statistical Effect size
Effect size is a quantitative measure of the magnitude of a phenomenon. It provides an indication of the practical significance of a result, independent of sample size, making it a critical complement to p-values in hypothesis testing. While p-values indicate whether a result is statistically significant, effect size reveals how strong the effect is, giving a clearer sense of its importance in real-world terms (Sullivan and Feinn 2012).
In the context of a categorical classification modeling challenge, effect size can serve as a critical pre-training tools to compare differences in feature space between classes.
Cohen’s d is commonly used for continuous feature variables, while odds ratios (ORs) are suitable for categorical variables. Both measures help quantify the magnitude of differences across populations.
The formula is
$d=\frac{\mu_{1}-\mu_{2}}{SD_{\text{pooled}}}$ (1)
where
  • $\mu_{1}$ and $\mu_{2}$ are the means of the two populations.
  • The pooled standard deviation ($SD_{\text{pooled}}$) is calculated as
    $SD_{\text{pooled}}=\sqrt{\frac{(n_{1}-1)\cdot SD_{1}^{2}+(n_{2}-1)\cdot SD_{2}^{2}}{n_{1}+n_{2}-2}}.$ (2)
The odds ratio (OR) measures the strength of association between a binary categorical variable (e.g., presence/absence of a condition) and group membership. The formula is
$OR=\frac{\text{Odds in Group 1}}{\text{Odds in Group 2}}.$ (3)
Odds are calculated as the ratio of the probability of an event occurring to the probability of it not occurring:
$\text{Odds}=\frac{p}{1-p}$ (4)
where $p$ is the probability of the event.
This allows us to calculate the individual effect sizes for each feature against a class (Hae-Young 2015). The cumulative effect size of an entire dataset against a class, can then be averaged for a resulting cumulative effect size score.
Hypothesis
This research attempts to ascertain whether or not certain descriptive statistical features of a dataset can be indicative of prospective model performance, and can provide a heuristic for data volume required to achieve certain generalizable performance benchmarks.
Corollary: The magnitude of Feature Effect size, and data volume, both impact model performance and generalizability, with a higher effect-size offsetting the need for large datasets, while a smaller effect size generally requiring a greater volume of data to achieve similar results.
Corollary: There is an upper-bound to performance irrespective of data size, and possibly effect size (i.e., as data size goes to infinite model performance plateaus).
The goal of this research: If a researcher has advanced knowledge the size of the dataset employable for training, and can calculate the effect size of features with respect to the labels, one can make a reasonable assumption on projected model performance.
To best explore this overall objective, we conducted two specifics experiments to determine whether or not such a relationship exists.
Hypothesis 1: There is a correlation between statistical effect-size and model performance up to a point.
Hypothesis 2: There is a correlation between statistical effect-size, data size and model performance up to a point
These trends are universal and applicable across datasets.
Methods
For this study, a dataset containing a sampling of adult census data was employed. The dataset was sourced from the UCI Machine Learning Repository and includes 48,842 observations across 14 features.
The following were the features along with data types:
Table 1: Features in the Census dataset.
Feature Name Description Type
Age Age of the individual Numerical
Workclass Type of employment Categorical
Fnlwgt Final weight; a census measure Numerical
Education Highest level of education completed Categorical
Education-Num Number of years of education Numerical
Marital-Status Marital status Categorical
Occupation Type of occupation Categorical
Relationship Relationship to head of household Categorical
Race Race of the individual Categorical
Sex Gender of the individual Categorical
Capital-Gain Income from investment Numerical
Capital-Loss Losses from investment Numerical
Hours-Per-Week Average hours worked per week Numerical
Native-Country Country of origin Categorical
Income Income level (target variable) Categorical
Opting for a binary classification problem, the label for this study was income, which was segmented into a binary categorical segmented as either greater than 50k or less than or equal to 50k per year (Becker and Kohavi 1996).
Standard preprocessing of the data was undertaken which included removing rows with null values, encoding categorical variables to ensure compatibility with specific models, and normalized numerical features. Categorical features were encoded using LabelEncoder, converting each category into a numerical value. All numerical features were standardized using  StandardScaler. The standardized numerical features and encoded categorical features are combined into a single feature matrix, which is then used for training and evaluating the machine learning models. Following the preprocessing step, a total of 33,000 samples remained.
The following set of experiments was then performed:
Experiment 1:
Determine if Effect Size Correlates to Performance Across datasets and feature mixes.
Experiment 1.1: Different subsets of data, with the same set of Features, (i.e., different mix of rows within the same dataset)
For this experiment the dataset was segmented into 66 subsets, each with 500 rows. All features remained identical across subsets, with only individual rows varying across datasets.
Experiment 1.1.1: Model performance was compared across the same set of features/different data.
For this experiment, the ”Income Greater Than 50k” binary categorical feature served as label, with the rest of the features serving as the feature set. For each data subset, the average effect size was calculated for the binary label we will aim to classify.
The effect size for numerical features was calculated using Cohen’s d, which measures the standardized difference between the means of two groups (e.g., individuals with income greater than \$50K and those with less than or equal to \$50K). For categorical features, the effect size is calculated using the Odds Ratio derived from a contingency table that compares the frequency distribution of the categories across the target variable (income).
After calculating the effect size for each individual feature, the average effect size across all features within a subset is computed. This average provides a single summary metric representing the overall effect size of the dataset in that particular subset.
An identical sequence of classifiers was trained on the models, employing different classification techniques. This was done to try and generalize the approach across learning model paradigms. Specifically, for every experiment, we employed the same four machine-learning models:
  1. 1.
    Logistic Regression
  2. 2.
    Decision Tree
  3. 3.
    Random Forest
  4. 4.
    Neural Network.
This yielded 66 data points of model performance vs. averaged effect size, allowing us to compare those two elements by calculating a direct correlation.
Experiment 1.1.2: Attempted to employ a Different Feature mix on the same rows, (i.e., utilizing the same hundred subsets of data but with a different set of features and labels.) Specifically the income label was reintroduced into the feature-mix and this time the ”sex” binary feature was employed as the label. The same set of models were run and performance metrics calculated, yielding another 66 datapoints.
Experiment 1.2: Compare model performance across a mix features with same underlying data
Experiment 1.2.1: For this experiment, model performance across a differing set of features with the same underlying data was compared. In this experiment the ”income greater than 50k” binary was retained as a label, while the rest of the features were retained as a feature set. For this experiment, a series of subsets were devised. For each subset one of the features was dropped, and then the averaged effect size for the binary label was recalculated on the reduced feature space. An identical sequence of classifiers was then trained, and model performance and associated averaged effect size were calculated.
This produced 14 data points of model performance vs. effect size.
Experiment 1.2.2: In a parallel fashion to experiment 1.1.2, we reran 1.2.1, but this time used the ”sex” binary label as a feature. An additional 14 data points were generated.
Experiment 2:
Compare the Relationship Between Effect Size and Learning Curve Slope.
This experiment sought to determine if there a relationship between the slope of a learning curve (the rate at which a model’s error rate decreases with respect to training data size) and the effect size, irrespective of overall model performance. The hypothesis being tested was whether or not a more dramatic effect size would indicate a cleaner dataset, requiring fewer samples for a model to ‘bottom out’ in terms of error. The critical distinction here was that it would not necessarily predict model performance, but rather how large a dataset was required for convergence on an optimal model.
For each model in experiment 1, where a previously calculated effect size was obtained, a learning curve/error rate for both training and validation sets was plotted.
Experiment 2.1: Correlation of Effect Size to Slope of Learning Curve
For each model, an associated learning curve on a validation set was plotted. This model generally adhered to a logrithmic function with an associated coefficient, beginning elevated, but converging on an optimal error rate with the addition of more training data. This slope was then plotted against the associated Effect Size and a correlation computed.
Experiment 2.2: Correlation of Effect Size to Ratio of Error between Training and Validation Sets
For each model, in addition to learning curve on the validation set, a learning curve on the training set was also plotted. These models tended to start with virtually no error, as a small amount of training data allowed for significant overfitting. As more training data is introduced, the error rate increases until it converges with the validation set’s error rate. Pairing both sets of plots, for each point, the magnitude of the difference in error between the training and validation sets was calculated. This resulted in a linearly decreasing amount as error rates converge. Then take slope of this linear line representing the magnitude of error difference was then extracted and correlated against Effect Size.
Results
Figure 1: Experiment 1.1.1 (income as label) with $R^{2}=0.0132$.
Figure 2: Experiment 1.1.2 (gender as label) with $R^{2}=0.0011$.
Figure 3: Experiment 1.2.1 (income as label) with $R^{2}=0.0142$.
Figure 4: Experiment 1.2.2 (gender as label) with $R^{2}=0.091$.
Experiment 1
Table 2: Summary of results for Experiment 1.
Exp. No. Experiment Name R2 Value
1.1.1 Income label, identical features, different subsets 0.0132
1.1.2 Sex label, identical features, different subsets 0.0011
1.2.1 Income label, different feature mix 0.0142
1.2.2 Sex label, different feature mix 0.091
Table 2 details the experimental results.
Overall the results demonstrably show that effect size has essentially no predictive value on the resulting model’s performance. As demonstrated by the corresponding R2 values, virtually none of the variance in model performance is attributable to statistical effect size.
While this work does not preclude the possibility of alternative statistical measures possibly serving as a metric for resulting model performance, thus far, such pre-hoc assessments remain elusive.
As ML algorithms are designed to maximize the performance of their outputs and minimize error irrespective of the input data, effective learning mechanisms can employ nonlinear and nonstatistical approaches to modeling data that cannot be effectively determined pre-hoc using relatively straight-forward linear statistical measures. Ultimately, this experiment suggests that model performance may have to be determined by some other factor of the dataset along with the specific model utilized.
Figure 5: Experiment 2.1 with $R^{2}=0.0054$.
Figure 6: Experiment 2.2 with $R^{2}=0.0007$.
Experiment 2 again fairly dramatically demonstrates that there is almost no correlation between effect size and data sufficiency. Experiment 2.1 suggests there is no correlation between the effect size and slope of the validation learning curve. Experiment 2.2 indicates no correlation between effect size and slope of the magnitude of the difference in error between the training and validation sets. The data points and R-Squared values suggest that determining a sufficient sample size to train your data is dependent on some other factor of our dataset along with the specific model utilized.
Discussion
Although given the limited nature of the experiments reported here, we cannot exhaustively conclude that there is no easily computable pre-hoc measure for data-sufficiency and model efficacy, it would not be surprising if this ultimately proved to be the case. Afterall, the nature of Machine Learning discards statistical certainty for predictive performance, thereby allowing for greater predictive inference than mere statistical modeling would otherwise allow. It would therefore be more surprising were such a metric easily obtainable.
Nevertheless, we believe the overriding goal of an effective pre-hoc assessment is still possible, albeit with a little bit more circuitous work needed to arrive at it.
In parallel with this work, we have conducted experiments on cross-model analysis, and have submitted for publication initial results that suggest that when comparing the internal activations of neural networks under varying conditions, models with similar behavioral architectures tend to exhibit similar levels of robustness. This suggests that if we train a model that is highly correlated to another well-established model, it will exhibit similar levels of generalizability (Oskouie, Levine, and Sarrafzadeh 2024).
This offers a potential next step in our own efforts here. If we can, from the outset, determine the likely structure of a model trained on a dataset will likely result in, and then correlate this prospective model to a preexisting one with known robustness, this may yield a useful metric on the likely robustness a model trained on a certain dataset is likely to exhibit prior to training it.
Conclusion
In this work, we attempted to demonstrate whether or not discrete statistical analyses of a dataset, in this case the Effect Size demonstrating the relative magnitude of dissonance among feature values across categorical values, could provide a useful indication of a resulting ML model’s performance or the sufficiency of the dataset size to effectively converge on an optimal model. Our results indicate that at least this particular metric bares little correlation to resulting model performance or data sufficiency. Nevertheless, we explore potential alternatives to advance the stated goal of better understanding the sufficiency of a dataset to produce a robust ML model prior to undertaking training and validation.

### v3

@@
 Our results appear to indicate that this is not an effective heuristic for determining adequate sample size or projecting model performance, and therefore that additional work is still needed to better prospectively assess adequacy of data.
-Codehttps://colab.research.google.com/drive/17cZ8pWYAQzCcENLgoUS01cRj5uLWi8ta?usp=sharing
-Datasetshttp://archive.ics.uci.edu/dataset/2/adult
-Introduction
+  1. 1.
+  2. 2.
+-A Introduction
 The sufficiency of a dataset, both in terms of size and representativeness, is critical to training an effective Machine Learning model.
-Failure to train on data of adequate quality is likely to result in models that dramatically under-perform in production environments (Friedland 2024).
+Failure to train on data of adequate quality is likely to result in models that dramatically under-perform in production environments [7].
 Training machine learning models on inadequate data presents several challenges that can significantly impact model performance and generalizability.
 While the problem of ’inadequacy’ of data is a multifaceted one, one significant component of it is a lack of sufficient data volume, which in-turn likely impacts the representativeness of the data sample towards the overall population undermining the ability of a model to learn meaningful patterns, resulting in overfitting, where the model performs well on the training data but struggles to generalize to unseen data leading to poor outcomes when deployed in real-world scenarios.
-This owes to the fact that if the dataset used for training does not accurately reflect the conditions and distributions the model will encounter in real-world applications, the model will struggle to generalize beyond the narrow scope of its training (Masiha et al.
-2021).
+This owes to the fact that if the dataset used for training does not accurately reflect the conditions and distributions the model will encounter in real-world applications, the model will struggle to generalize beyond the narrow scope of its training [10].
 Consequently, the model’s assumptions about the domain may be incorrect, leading to suboptimal decision-making in practical applications.
 In cases where the dataset is particularly small or inadequate, machine learning models—especially complex ones like deep learning architectures—fail to learn the intricate relationships within the data.
-Complex models require large amounts of data to effectively capture subtle patterns (Keshari et al.
-2020).
+Complex models require large amounts of data to effectively capture subtle patterns [9].
 Without sufficient data, the model may not converge, or it may underfit, meaning it performs poorly on both the training and test sets due to an inability to learn the underlying patterns.
 Moreover, the use of inadequate data increases variability in model performance.
-Small datasets are more sensitive to changes in initialization values, training procedures, or even small differences in the data samples used during training (Dodge et al.
-2020).
+Small datasets are more sensitive to changes in initialization values, training procedures, or even small differences in the data samples used during training [6].
 As a result, models trained on inadequate data exhibit inconsistent and unpredictable behavior, making them unreliable in different testing or production environments.
@@
 This owes both to the upfront challenges and associated costs of data collection, and with the advent of increasingly costly training runs, the investment costs associated with training.
-This is a widely recognized need across research domains, with many experimental protocols requiring power analyses, or equivalent statistical studies to prospectively determine the requisite amount of data collection required to run studies effectively.(Bausell and Li 2002)
-However, in spite of the obvious utility of such an assessment, it remains elusive for machine learning model developers.
+This is a widely recognized need across research domains, with many experimental protocols requiring power analyses, or equivalent statistical studies to prospectively determine the requisite amount of data collection required to run studies effectively.[2] However, in spite of the obvious utility of such an assessment, it remains elusive for machine learning model developers.
 As already noted, many methods for external dataset-based validation but are post-hoc model dependent.
-Background
-Prospective Analysis of Data Sufficiency
+-B Background
+-B1 Prospective Analysis of Data Sufficiency
 In most academic disciplines, conducting a sufficiently robust study design is arguably more important than the statistical analysis that succeeds it.
-After-all, a poorly designed study may be unsalvageable after the fact, whereas a poorly analyzed study can simply be re-analyzed once errors in methodology are determined (BMJ 2020).
+After-all, a poorly designed study may be unsalvageable after the fact, whereas a poorly analyzed study can simply be re-analyzed once errors in methodology are determined [5].
 A critical component in study design is the determination of the appropriate sample size.
@@
 Attempting to determine data sufficiency for machine learning (ML) models is a particularly vexing challenge given the nature of machine learning.
-Unlike statistics, ML does not attempt to asset any factual relationship between label and feature set (Bennett et al.
-2022); Rather ML models are almost mercenary in nature.
+Unlike statistics, ML does not attempt to asset any factual relationship between label and feature set [4]; Rather ML models are almost mercenary in nature.
 Charged with best accomplishing a given prediction or classification task, irrespective of any actual underlying statistical relationship.
 This relaxed focus allows for a more expansive inclusion of features for whom a statistically significant relationship to the label may not exist.
-Current approaches to determining sufficient sample size
+-B2 Current approaches to determining sufficient sample size
 There is a lack of approaches that exist for determining a sufficient sample size to train ML models.
-The absence of a robust pre-hoc approach that is broadly applicable remains a significant challenge in machine learning, largely stemming from the limited research conducted in this area (Balki et al.
-2019).
+The absence of a robust pre-hoc approach that is broadly applicable remains a significant challenge in machine learning, largely stemming from the limited research conducted in this area [1].
 The most common current approach is empirical analysis.
@@
 There are mathematical equations like those of Baum and Hausler that allow one to determine a sufficient number of samples through this model-based framework.
-However, this framework is not optimal for all tasks and has its trade-offs.(Balki et al.
-2019)
-Statistical Effect size
+However, this framework is not optimal for all tasks and has its trade-offs.[1]
+-B3 Statistical Effect size
 Effect size is a quantitative measure of the magnitude of a phenomenon.
 It provides an indication of the practical significance of a result, independent of sample size, making it a critical complement to p-values in hypothesis testing.
-While p-values indicate whether a result is statistically significant, effect size reveals how strong the effect is, giving a clearer sense of its importance in real-world terms (Sullivan and Feinn 2012).
+While p-values indicate whether a result is statistically significant, effect size reveals how strong the effect is, giving a clearer sense of its importance in real-world terms [11].
 In the context of a categorical classification modeling challenge, effect size can serve as a critical pre-training tools to compare differences in feature space between classes.
@@
 Both measures help quantify the magnitude of differences across populations.
-The formula is
-$d=\frac{\mu_{1}-\mu_{2}}{SD_{\text{pooled}}}$ (1)
-where
-  • $\mu_{1}$ and $\mu_{2}$ are the means of the two populations.
-  • The pooled standard deviation ($SD_{\text{pooled}}$) is calculated as
+The formula is $d=\frac{\mu_{1}-\mu_{2}}{SD_{\text{pooled}}},$ where $\mu_{1}$ and $\mu_{2}$ are the means of the two populations, and the pooled standard deviation ($SD_{\text{pooled}}$) is calculated as
 $SD_{\text{pooled}}=\sqrt{\frac{(n_{1}-1)\cdot SD_{1}^{2}+(n_{2}-1)\cdot SD_{2}^{2}}{n_{1}+n_{2}-2}}.$ (1)
@@
 $OR=\frac{\text{Odds in Group 1}}{\text{Odds in Group 2}}.$ (2)
-Odds are calculated as the ratio of the probability of an event occurring to the probability of it not occurring:
-$\text{Odds}=\frac{p}{1-p}$ (4)
-where $p$ is the probability of the event.
-This allows us to calculate the individual effect sizes for each feature against a class (Hae-Young 2015).
+Odds are calculated as the ratio of the probability of an event occurring to the probability of it not occurring $\text{Odds}=\frac{p}{1-p},$ where $p$ is the probability of the event.
+This allows us to calculate the individual effect sizes for each feature against a class [8].
 The cumulative effect size of an entire dataset against a class, can then be averaged for a resulting cumulative effect size score.
-Hypothesis
+-C Hypothesis
 This research attempts to ascertain whether or not certain descriptive statistical features of a dataset can be indicative of prospective model performance, and can provide a heuristic for data volume required to achieve certain generalizable performance benchmarks.
@@
 Corollary: There is an upper-bound to performance irrespective of data size, and possibly effect size (i.e., as data size goes to infinite model performance plateaus).
+Figure 1: Experiment 1.1.1 (income as label): Mean absolute effect size plotted against model performance (accuracy, precision, recall, and F1-score).
+Each point represents a model (distinguished by color) trained on a distinct subset.
 The goal of this research: If a researcher has advanced knowledge the size of the dataset employable for training, and can calculate the effect size of features with respect to the labels, one can make a reasonable assumption on projected model performance.
@@
 Hypothesis 1: There is a correlation between statistical effect-size and model performance up to a point.
-Hypothesis 2: There is a correlation between statistical effect-size, data size and model performance up to a point
+Hypothesis 2: There is a correlation between statistical effect-size, data size, and model performance up to a point.
 These trends are universal and applicable across datasets.
-Methods
-For this study, a dataset containing a sampling of adult census data was employed. The dataset was sourced from the UCI Machine Learning Repository and includes 48,842 observations across 14 features.
-The following were the features along with data types:
-Table 1: Features in the Census dataset.
+I Methods
+For this study, a dataset containing a sampling of adult census data was employed.
+The dataset was sourced from the UCI Machine Learning Repository and includes 48,842 observations across 14 features.
+The features along with data types are shown at Table I.
+TABLE I: Features in the Census dataset.
 Feature Name Description Type
@@
 Income Income level (target variable) Categorical
-Opting for a binary classification problem, the label for this study was income, which was segmented into a binary categorical segmented as either greater than 50k or less than or equal to 50k per year (Becker and Kohavi 1996).
+Figure 2: Experiment 1.1.2 (gender as label): Mean absolute effect size plotted against model performance (accuracy, precision, recall, and F1-score).
+Each point represents a model (distinguished by color) trained on a distinct subset.
+Figure 3: Experiment 1.2.1 (income as label): Effect size of removed feature plotted against model performance drop (accuracy drop, precision drop, recall drop, and F1-score drop).
+Each point represents a model (distinguished by color) with a different feature removed.
+Figure 4: Experiment 1.2.2 (gender as label): Effect size of removed feature plotted against model performance drop (accuracy drop, precision drop, recall drop, and F1-score drop).
+Each point represents a model (distinguished by color) with a different feature removed.
+Figure 5: Experiment 2.1 (left): Mean absolute effect size plotted against slope of the learning curve.
+Experiment 2.2 (right): Mean absolute effect size plotted against slope of the difference between training learning curve and validation learning curve.
+Both experiments set income as label.
+Each point represents a model (distinguished by color) trained on a distinct subset.
+Opting for a binary classification problem, the label for this study was income, which was segmented into a binary categorical segmented as either greater than 50k or less than or equal to 50k per year [3].
 Standard preprocessing of the data was undertaken which included removing rows with null values, encoding categorical variables to ensure compatibility with specific models, and normalized numerical features.
@@
 The following set of experiments was then performed:
-Experiment 1:
+I-1 Experiment 1
 Determine if Effect Size Correlates to Performance Across datasets and feature mixes.
@@
 This was done to try and generalize the approach across learning model paradigms.
-Specifically, for every experiment, we employed the same four machine-learning models:
-  1. 1.
-    Logistic Regression
-  2. 2.
-    Decision Tree
-  3. 3.
-    Random Forest
-  4. 4.
-    Neural Network.
-This yielded 66 data points of model performance vs. averaged effect size, allowing us to compare those two elements by calculating a direct correlation.
+Specifically, for every experiment, we employed the same four machine-learning models: Logistic Regression, Decision Tree, Random Forest, and Neural Network.
+This yielded 66 data points of model performance vs.
+averaged effect size, allowing us to compare those two elements by calculating a direct correlation.
 Experiment 1.1.2: Attempted to employ a Different Feature mix on the same rows, (i.e., utilizing the same hundred subsets of data but with a different set of features and labels.) Specifically the income label was reintroduced into the feature-mix and this time the ”sex” binary feature was employed as the label.
@@
 In this experiment the ”income greater than 50k” binary was retained as a label, while the rest of the features were retained as a feature set.
+A baseline on all performance metrics was initially calculated.
 For this experiment, a series of subsets were devised.
-For each subset one of the features was dropped, and then the averaged effect size for the binary label was recalculated on the reduced feature space.
-An identical sequence of classifiers was then trained, and model performance and associated averaged effect size were calculated.
-This produced 14 data points of model performance vs. effect size.
+For each subset one of the features was dropped, and the effect size of the dropped feature was calculated.
+An identical sequence of classifiers was then trained, and model performance drop was calculated.
+This produced 14 data points of model performance drop vs.
+effect size of removed feature.
 Experiment 1.2.2: In a parallel fashion to experiment 1.1.2, we reran 1.2.1, but this time used the ”sex” binary label as a feature. An additional 14 data points were generated.
-Experiment 2:
+I-2 Experiment 2
 Compare the Relationship Between Effect Size and Learning Curve Slope.
@@
 Then take slope of this linear line representing the magnitude of error difference was then extracted and correlated against Effect Size.
-Results
-Figure 1: Experiment 1.1.1 (income as label) with $R^{2}=0.0132$.
-Figure 2: Experiment 1.1.2 (gender as label) with $R^{2}=0.0011$.
-Figure 3: Experiment 1.2.1 (income as label) with $R^{2}=0.0142$.
-Figure 4: Experiment 1.2.2 (gender as label) with $R^{2}=0.091$.
-Experiment 1
-Table 2: Summary of results for Experiment 1.
-Exp. No. Experiment Name R2 Value
-1.1.1 Income label, identical features, different subsets 0.0132
-1.1.2 Sex label, identical features, different subsets 0.0011
-1.2.1 Income label, different feature mix 0.0142
-1.2.2 Sex label, different feature mix 0.091
-Table 2 details the experimental results.
-Overall the results demonstrably show that effect size has essentially no predictive value on the resulting model’s performance.
-As demonstrated by the corresponding R2 values, virtually none of the variance in model performance is attributable to statistical effect size.
-While this work does not preclude the possibility of alternative statistical measures possibly serving as a metric for resulting model performance, thus far, such pre-hoc assessments remain elusive.
-As ML algorithms are designed to maximize the performance of their outputs and minimize error irrespective of the input data, effective learning mechanisms can employ nonlinear and nonstatistical approaches to modeling data that cannot be effectively determined pre-hoc using relatively straight-forward linear statistical measures.
-Ultimately, this experiment suggests that model performance may have to be determined by some other factor of the dataset along with the specific model utilized.
-Figure 5: Experiment 2.1 with $R^{2}=0.0054$.
-Figure 6: Experiment 2.2 with $R^{2}=0.0007$.
-Experiment 2 again fairly dramatically demonstrates that there is almost no correlation between effect size and data sufficiency.
-Experiment 2.1 suggests there is no correlation between the effect size and slope of the validation learning curve.
-Experiment 2.2 indicates no correlation between effect size and slope of the magnitude of the difference in error between the training and validation sets.
-The data points and R-Squared values suggest that determining a sufficient sample size to train your data is dependent on some other factor of our dataset along with the specific model utilized.
-Discussion
-Although given the limited nature of the experiments reported here, we cannot exhaustively conclude that there is no easily computable pre-hoc measure for data-sufficiency and model efficacy, it would not be surprising if this ultimately proved to be the case.
-Afterall, the nature of Machine Learning discards statistical certainty for predictive performance, thereby allowing for greater predictive inference than mere statistical modeling would otherwise allow.
-It would therefore be more surprising were such a metric easily obtainable.
-Nevertheless, we believe the overriding goal of an effective pre-hoc assessment is still possible, albeit with a little bit more circuitous work needed to arrive at it.
-In parallel with this work, we have conducted experiments on cross-model analysis, and have submitted for publication initial results that suggest that when comparing the internal activations of neural networks under varying conditions, models with similar behavioral architectures tend to exhibit similar levels of robustness.
-This suggests that if we train a model that is highly correlated to another well-established model, it will exhibit similar levels of generalizability (Oskouie, Levine, and Sarrafzadeh 2024).
-This offers a potential next step in our own efforts here.
-If we can, from the outset, determine the likely structure of a model trained on a dataset will likely result in, and then correlate this prospective model to a preexisting one with known robustness, this may yield a useful metric on the likely robustness a model trained on a certain dataset is likely to exhibit prior to training it.
-Conclusion
-In this work, we attempted to demonstrate whether or not discrete statistical analyses of a dataset, in this case the Effect Size demonstrating the relative magnitude of dissonance among feature values across categorical values, could provide a useful indication of a resulting ML model’s performance or the sufficiency of the dataset size to effectively converge on an optimal model.
-Our results indicate that at least this particular metric bares little correlation to resulting model performance or data sufficiency.
-Nevertheless, we explore potential alternatives to advance the stated goal of better understanding the sufficiency of a dataset to produce a robust ML model prior to undertaking training and validation.
+II Results
+Multiple inferential metrics were utilized to evaluate results.
+Pearson correlation (r with associated p-value) to evaluate linear relationship.
+The r-squared value was derived to determine how much of the variance in the dependent variable can be explained by the independent variable.
+Spearman correlation ($\rho$ with associated p-value) to detect monotonic relationship.
+Kendall’s Tau ($\tau$ with associated p-value) to establish whether any consistent ordering between variables exists.
+A 95% bootstrap confidence interval for pearson and spearman was also implemented for further validation of results, since experiments were run on a single dataset.
+Here the data was resampled 1000 times to see if any correlation that does exist is sound based on if it shows up consistently.
+Table II details the experimental results and plots for each experiment can be found in Figures 1,2, 3, 4, and 5.
+TABLE II: Complete Summary of Inferential Metrics Across All Experiments
+Exp. Metric Pearson $r$ $p$ $R^{2}$ Spearman $\rho$ $p$ Kendall $\tau$ $p$ Slope Pearson CI Spearman CI
+Experiment 1.1.1 — Income Label, Identical Features
+1.1.1 Accuracy -0.083 0.181 0.007 -0.052 0.399 -0.035 0.417 -0.028 [-0.205, 0.047] [-0.172, 0.067]
+1.1.1 Precision -0.056 0.362 0.003 -0.033 0.590 -0.024 0.568 -0.059 [-0.164, 0.054] [-0.152, 0.088]
+1.1.1 Recall 0.101 0.102 0.010 0.121 0.049 0.080 0.057 0.101 [-0.013, 0.217] [0.007, 0.241]
+1.1.1 F1 Score 0.061 0.327 0.004 0.078 0.204 0.049 0.236 0.051 [-0.058, 0.171] [-0.038, 0.198]
+Experiment 1.1.2 — Sex Label, Identical Features
+1.1.2 Accuracy -0.173 0.005 0.030 -0.150 0.014 -0.101 0.018 -0.072 [-0.287, -0.056] [-0.272, -0.038]
+1.1.2 Precision -0.139 0.024 0.019 -0.123 0.046 -0.084 0.045 -0.051 [-0.254, -0.014] [-0.238, 0.001]
+1.1.2 Recall -0.145 0.018 0.021 -0.086 0.165 -0.057 0.169 -0.065 [-0.265, -0.027] [-0.204, 0.025]
+1.1.2 F1 Score -0.185 0.003 0.034 -0.144 0.019 -0.097 0.021 -0.058 [-0.305, -0.066] [-0.263, -0.029]
+Experiment 1.2.1 — Income Label, Feature Removal
+1.2.1 Accuracy Drop -0.123 0.365 0.015 -0.003 0.982 -0.017 0.864 -0.001 [-0.387, 0.164] [-0.263, 0.275]
+1.2.1 Precision Drop -0.138 0.312 0.019 0.012 0.930 -0.009 0.926 -0.004 [-0.402, 0.157] [-0.254, 0.285]
+1.2.1 Recall Drop -0.118 0.387 0.014 -0.055 0.688 -0.043 0.666 -0.004 [-0.373, 0.111] [-0.326, 0.232]
+1.2.1 F1 Drop -0.141 0.302 0.020 -0.039 0.774 -0.031 0.744 -0.004 [-0.406, 0.143] [-0.330, 0.249]
+Experiment 1.2.2 — Sex Label, Feature Removal
+1.2.2 Accuracy Drop -0.068 0.621 0.005 -0.154 0.258 -0.105 0.286 -0.001 [-0.344, 0.202] [-0.429, 0.132]
+1.2.2 Precision Drop -0.026 0.850 0.001 -0.141 0.301 -0.093 0.328 -0.000 [-0.207, 0.204] [-0.397, 0.135]
+1.2.2 Recall Drop -0.053 0.700 0.003 -0.055 0.690 -0.038 0.702 -0.000 [-0.289, 0.192] [-0.318, 0.221]
+1.2.2 F1 Drop -0.071 0.606 0.005 -0.150 0.271 -0.097 0.307 -0.000 [-0.368, 0.220] [-0.415, 0.142]
+Experiment 2.1 — Learning Curve Slope
+2.1 LC Slope -0.115 0.099 0.013 -0.185 0.007 -0.129 0.006 -0.042 [-0.274, 0.049] [-0.326, -0.031]
+Experiment 2.2 — Train–Validation Gap Slope
+2.2 Gap Slope 0.065 0.349 0.004 0.103 0.138 0.072 0.124 0.000 [-0.097, 0.206] [-0.049, 0.236]
+II-A Experiment 1
+Experiment 1.1.1: The pearson r-values do not indicate any relationship and the corresponding p-values indicate that any relationship that may exist is statistically insignificant.
+Figure 1 is consistent with this result and shows no linear trend in the data.
+The spearman-rho value for recall suggests a very weak positive monotonic relationship.
+The corresponding p-value indicates that although the relationship is weak, it is likely not a random result.
+The spearman for the other performance metrics does not reveal any statistically detectable relationships.
+The slope for recall is approximately 10% and the spearman confidence interval for recall excludes zero.
+These results align with the positive monotic relationship observed and confirm consistency.
+Experiment 1.1.2: The pearson r-values and corresponding p-values indicate that a linear negative relationship is statistically detectable, but is very weak.
+As the absolute average effect size increases the performance seems to slightly decrease.
+We see in Accuracy and F1-Score this relationship is slightly stronger.
+The spearman-$\rho$ values for Accuracy, Precision, and F1-Score suggest a very weak negative monotonic relationship.
+And the corresponding p-values for these indicate that although the relationship is weak, it is likely not a random result.
+The kendall-tau and slope values are quite small, but they do align with the negative trend observed.
+In the confidence intervals, the pearson excludes zero for all performance metrics.
+Further confirming that although the relationship is weak, it is consistent.
+Similarly with Spearman confidence intervals, zero is excluded for Accuracy and F1-Score.
+Figure 2 visually reflects these results as there is a very slight downtrend when compared to to Figure 1.
+Experiment 1.2.1: The pearson r-values suggest a weak negative correlation, but corresponding p-values are very high.
+This indicates that statistically we cannot reject the null hypothesis.
+Figure 3 and the remaining metrics do not suggest any statistically detectable relationships between the effect size of the removed feature and the change in performance.
+Experiment 1.2.2 The pearson r-values and kendall values show no relationship.
+The spearman $\rho$-values for accuracy drop, precision drop, and F1 drop signal a weak negative monotonic relationship.
+However, the corresponding p-values for these three inferential metrics confirm that they are statistically insignificant.
+Figure 4 and the confidence intervals align with these results, showing no detectable consistent relationship.
+II-B Experiment 2
+Experiment 2.1: The pearson r-value for the experiment indicates a negative correlation, but the p-value considers this statistically insignficant.
+The spearman $\rho$-value suggests a weak negative montonic relationship between absolute average effect size and the slope of the learning curve.
+The corresponding p-value confirms this relationship is unlikely due to random chance.
+The kendall value and its p-value support this observation and suggest a very weak relationship as well.
+As absolute average effect size goes up, the slope of the learning curve slightly decreases.
+The spearman confidence interval confirms consistency.
+Experiment 2.2 This experiment indicates no relationship between effect size and slope of the magnitude of the difference in error between the training and validation sets.
+The data points and metrics suggest that determining a sufficient sample size to train your data is dependent on some other factor of our dataset along with the specific model utilized.
+Figure 5 is consistent with the results of both experiments.
+On the left panel, there is a slight downtrend, though it is not prominent.
+The right panel reveals no observable relationship.
+III Discussion
+Our findings suggest that simple, univariate summaries of class separability—even when aggregated across features via mean absolute effect size—are not reliable proxies for downstream model quality or sample-size sufficiency in realistic, mixed-type tabular data.
+A key reason is that the “average effect size” statistic compresses a high-dimensional, heterogeneous feature set into a single scalar, discarding the structure that learning algorithms exploit.
+In the Adult dataset, many informative features are (i) nonlinearly expressed (e.g., interactions between education, occupation, capital gains, and hours worked) and (ii) embedded in multi-category variables with class-conditional distributions that are not well summarized by a single odds ratio.
+As a result, a dataset may look weakly separable when features are examined one at a time, yet still be easy to learn because of features interactions; alternatively, it may look strongly separable due to redundant or spurious correlations that do not generalize.
+Moreover, the mapping from data to performance is model- and metric-dependent.
+Effect size is a population statistic independent of the learner, whereas generalization error depends on inductive bias, regularization, optimization, and the evaluation metric.
+For example, tree-based models can exploit nonlinear threshold effects and high-order feature interactions, while linear models rely more directly on marginal separability under the chosen feature representation.
+In this setting, a single effect-size aggregate cannot capture whether the available feature is aligned with the hypothesis class of the model family.
+This helps explain why correlations were small even when statistically detectable.
+Taken together, our study indicates that prospective data sufficiency assessment likely requires statistics that (i) respect feature interactions and representation, and (ii) incorporate at least minimal model assumptions.
+Two practical directions are promising.
+First, replace univariate effect-size aggregation with multivariate measures of class separability that are sensitive to interactions.
+Second, adopt model-aware pre-hoc predictors that are still inexpensive relative to full training.
+These approaches preserve the goal of prospective assessment while acknowledging that “data quality” is inseparable from the learner and representation used.
+More broadly, our results suggest that while simple descriptive statistics are useful for understanding marginal feature behavior, they are not reliable stand-alone predictors of model performance or sample-size sufficiency.
+IV Conclusion
+This paper investigated whether a simple, pre-training statistic—mean absolute feature effect size can serve as a practical heuristic for (i) predicting downstream classification performance and (ii) anticipating sample-size sufficiency via learning-curve convergence behavior on mixed-type tabular data.
+Across two complementary settings (varying row subsets with fixed features and varying feature sets via leave-one-feature-out) and four model families (logistic regression, decision tree, random forest, and a neural network), the observed associations between aggregated effect size and predictive metrics were consistently weak and often inconsistent.
+These results suggest that effect-size aggregation should not be viewed as a stand-alone predictor of model success or dataset adequacy.
+Future work toward data sufficiency assessment should therefore emphasize multivariate measures that preserve interaction information and low-cost, model-aware predictors that maintain practicality while improving decision relevance for experimental design and data acquisition.